OpenStudy (anonymous):

Factor this expression as a difference of squares: 4-(2a+3b)^2

Parth (parthkohli):

\[2^2 - (2a + 3b)^2\]Do you know the formula?

OpenStudy (anonymous):

Well, I haven't seen problems like this a years, I just need a refresher on how to understand it properly. It just says what it says above.

jimthompson5910 (jim_thompson5910):

use the difference of squares rule x^2 - y^2 = (x-y)(x+y)

OpenStudy (anonymous):

use the basic formula for the diff f squares : \[a^2-b^2=(a-b)(a+b)\] and then substitute a for 2 and b for the expression in the bracket

OpenStudy (anonymous):

you will then get : \[2^2-(2a+3b)^2 = (2-[2a+3b])(2+[2a+3b])\] \[= (2-2a-3b)(2+2a+3b)\]

OpenStudy (anonymous):

Can you continue from there?